Hello,
I am trying to solve this problem, I am doing radicals and rational exponents and I am having problems solving to questions that needs to be done by tomorrow morning could someone please help with these two problems? First problem is 25^-3/2 and the second problem is (-27x^9)^1/3.
Thanks very much for your help!

Answer by edjones(8007)

25^(-3/2)
=1/25^(3/2)
=1/15625^(1/2)
=1/125
.
Ed